OpenStudy (anonymous):

How would I solve 125^(-1/3)?

OpenStudy (kropot72):

You can use the following" \[n ^{-\frac{1}{a}}=\frac{1}{n ^{\frac{1}{a}}}=\frac{1}{\sqrt[a]{n}}\]

OpenStudy (anonymous):

125^(1/3) = (5^3)^(1/3) .............. 125 = 5^3 = 5 ........... (3 and 1/3 will cancel out) Hope this helps @DanielAcedo